OpenStudy (anonymous):

Solve the equation: \[\frac{ t }{ 3}-\frac{ 1 }{ 2}=1\]

OpenStudy (anonymous):

first you are going to need a common denominator. 6 will work nicely 6((t/3)-(1/2))=1*6 2t-3=6 then add three to both sides 2t=9 divide by 2 t= 9/2

OpenStudy (anonymous):

A. 4.1 B. 5.0 C. 4.2 D. 4.5

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

no, it would be d because 9/2 is 4.5

OpenStudy (anonymous):

thank you did you divide 9/2? thats how you got it ?

OpenStudy (anonymous):

i did get 4.5 by dividing 9 by 2, yes